OpenStudy (anonymous):

verify the expression. 4sinθ/(1-sin^2⁡〖ϴ 〗 )=2 sec^2⁡ϴ

OpenStudy (loser66):

recheck the question, please

OpenStudy (mathmath333):

can u write in with equation editor

OpenStudy (anonymous):

the initial question was 1/(1-sinϴ)+1/(1+sinϴ)=2 sec^2⁡ϴ

OpenStudy (anonymous):

the ?s are supposed to be theta

OpenStudy (mathmath333):

ok wait

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \dfrac{1}{(1-sin\theta)}+\dfrac{1}{(1+sin\theta)}\\~\\ =\dfrac{1+sin\theta+1-sin\theta}{(1-sin^2\theta)}\\~\\ =\dfrac{2}{cos^2\theta}\\~\\ =2sec^2\theta}\end{align}\)

OpenStudy (anonymous):

wow what the hell was i doing lol. thank you!

OpenStudy (mathmath333):

lol